# OneRanker: Unified Generation and Ranking with One Model in Industrial Advertising Recommendation

### 04 Mar 2026 v2

@@ 1. Introduction
Finally, to further enhance generation-ranking consistency and complete the unified framework, we propose dual-side consistency guarantees. On the input side, the ranking decoder constructs its Key/Value from both Step 1’s original multi-interest representations and Step 2’s refined representations, ensuring ranking decisions fully inherit generation process information. On the output side, we introduce a Distributional Consistency (DC) mechanism that propagates business value signals from Step 3’s ranker back to Step 2’s generator as soft labels. Through a joint loss function $\mathcal{L}_{\text{total}}=\alpha\mathcal{L}_{\text{MTP}}+\beta\mathcal{L}_{\text{rank}}+\gamma\mathcal{L}_{\text{consistency}}$, we achieve gradient-level collaborative optimization, enabling the generator to “anticipate” ranking preferences during training and fundamentally preventing objective drift between stages.
@@ 4.1.2. Baselines and Evaluation Metrics
To comprehensively evaluate recommendation accuracy and ranking quality, we adopt Hit Ratio (HR@K) as our primary evaluation metrics.
@@ 4.1.4. Results
As illustrated in Table 1, we evaluate the performance of different architectures across various Hit Ratio (HR) metrics. Firstly, both HSTU and GPR serve as competitive baselines for capturing user interests. Notably, the GPR model demonstrates a clear advantage over HSTU across all measured intervals; for instance, it improves the HR@1 and HR@5 from 0.1741 and 0.4648 to 0.1824 and 0.4935, respectively. The proposed OneRanker model achieves a significant breakthrough, substantially outperforming both baselines by a wide margin. Specifically, OneRanker reaches an HR@1 of 0.2639, which represents a 44.7% improvement over the GPR model. Similar dominant gains are observed in HR@5 (0.6213) and HR@15 (0.8894). These results effectively validate that our integrated architecture, with its task-specific enhancements and value-aware ranking, can more precisely align model outputs with complex recommendation objectives, yielding superior performance across the board.
@@
Model/Metric HR@1 HR@3 HR@5 HR@10 HR@15
HSTU 0.1741 0.3508 0.4648 0.6604 0.7953
GPR 0.1824 0.3703 0.4935 0.6957 0.8207
OneRanker 0.2639 0.4959 0.6213 0.7945 0.8894
@@ 4.2.1. Ablation Studies on Key Structural Components (RQ2)
Model/Method HR@1 HR@3 HR@5 HR@10 HR@15
OneRanker (Full) 0.2639 0.4959 0.6213 0.7945 0.8894
w/o DC Loss 0.2629 0.4925 0.6173 0.7905 0.8867
w/o S2 token injection 0.2610 0.4912 0.6161 0.7897 0.8868
with S3 ranker only 0.2601 0.4901 0.6157 0.7880 0.8852
OneRanker S2 0.2214 0.4257 0.5448 0.7291 0.8435
w/o Target 0.2080 0.4030 0.5203 0.7066 0.8276
w/o Target & MDA 0.2035 0.3922 0.5066 0.6949 0.8212
Experimental results indicate that each core module we designed brings significant improvements to model performance, and the ablation experiments further verify it, as is shown in Table 2. We primarily use HR@5 for comparison.
In Step 2, compared to the Base (OneRanker S2) HR@5 performance of 0.5448, the performance drops to 0.5203 (-4.5%) after removing Target information (w/o Target), which strongly proves the necessity of the Fake-Item-Token mechanism, prompting the model to evolve from static user representations to dynamic-aware item semantic distribution representations, thereby capturing users’ real-time intentions more accurately. Furthermore, removing both the Target information and the Multi-task Decoupled Architecture (w/o Target & MDA) results in an HR@5 decline to 0.5066 (-7.0%) compared to the OneRanker S2 base, underscoring the critical role of the decoupled architecture in mitigating optimization tension within the shared representation space, effectively preventing interference between business value modeling and user interest modeling.
In Step 3, using the OneRanker with S3 ranker only (HR@5: 0.6157) as the baseline, the model already demonstrates a leap-forward boost compared to S2-only models, proving that fine-grained target-aware mechanisms are fundamental to effective ranking. Building upon this, the introduction of S2 information injection (HR@5: 0.6161) ensures that the ranker reuses the generation logic to better calibrate its predictions within the ”generation context”. Furthermore, incorporating the Distributional Consistency ${L}_{\text{DC}}$ (HR@5: 0.6173) explicitly constrains the output distribution between stages, mitigating target shift through consistency regularization. Ultimately, the OneRanker (Full) model achieves the optimal HR@5 of 0.6213. This improvement stems from a collaborative effect: while S2 provides enriched feature injection to S3, the Distributional Consistency loss facilitates essential value-driven feedback from S3 back to S2. This closed-loop interaction between input consistency and output consistency allows the model to reach a higher ceiling of performance.

### 12 Mar 2026 v3

@@ 1. Introduction
Finally, to further enhance generation-ranking consistency and complete the unified framework, we propose dual-side consistency guarantees. On the input side, the ranking decoder constructs its Key/Value from both Step 1’s original multi-interest representations and Step 2’s refined representations, ensuring ranking decisions fully inherit generation process information. On the output side, we introduce a Distributional Consistency (DC) mechanism that propagates business value signals from Step 3’s ranker back to Step 2’s generator as soft labels. Through a joint loss function $\mathcal{L}_{\text{total}}=\alpha\mathcal{L}_{\text{MTP}}+\beta\mathcal{L}_{\text{rank}}+\gamma\mathcal{L}_{\text{DC}}$, we achieve gradient-level collaborative optimization, enabling the generator to “anticipate” ranking preferences during training and fundamentally preventing objective drift between stages.
@@ 4.1.2. Baselines and Evaluation Metrics
To comprehensively evaluate recommendation accuracy and ranking quality, we adopt Hit Ratio (HR@K) and Normalized Discounted Cumulative Gain (NDCG@K) as our primary evaluation metrics.
@@ 4.1.4. Results
As illustrated in Table 1, we evaluate the performance of different architectures across various Hit Ratio (HR) metrics. Firstly, both HSTU and GPR serve as competitive baselines for capturing user interests. Notably, the GPR model demonstrates a consistent advantage over HSTU across all measured intervals and metrics. For instance, GPR improves the HR@1 and NDCG@1 from 0.1741 and 0.6742 to 0.1824 and 0.6823, respectively. The proposed OneRanker model achieves a significant breakthrough, substantially outperforming both baselines by a wide margin. Specifically, OneRanker reaches an HR@1 of 0.2639, which represents a 44.7% relative improvement over the GPR model. Similar dominant gains are observed in ranking quality metrics; for example, NDCG@5 increases from 0.6818 to 0.7904, and NDCG@15 increases from 0.7445 to 0.8206. These results effectively validate that our architecture, with its task-specific enhancements and value-aware ranking, can more precisely align model outputs with complex recommendation objectives, yielding superior performance across the board.
@@
Model/Method HR@1 HR@3 HR@5 HR@10 HR@15 NDCG@1 NDCG@3 NDCG@5 NDCG@10 NDCG@15
HSTU 0.1741 0.3508 0.4648 0.6604 0.7953 0.6742 0.6712 0.6763 0.7025 0.7396
GPR 0.1824 0.3703 0.4935 0.6957 0.8207 0.6823 0.6776 0.6818 0.7070 0.7445
OneRanker 0.2639 0.4959 0.6213 0.7945 0.8894 0.8102 0.7954 0.7904 0.7970 0.8206
@@ 4.2.1. Ablation Studies on Key Structural Components (RQ2)
Model/Method HR@1 HR@3 HR@5 HR@10 HR@15 NDCG@1 NDCG@3 NDCG@5 NDCG@10 NDCG@15
OneRanker (Full) 0.2639 0.4959 0.6213 0.7945 0.8894 0.8102 0.7954 0.7904 0.7970 0.8206
w/o DC Loss 0.2629 0.4925 0.6173 0.7905 0.8867 0.8059 0.7912 0.7865 0.7936 0.8180
w/o S2 token injection 0.2610 0.4912 0.6161 0.7897 0.8868 0.8047 0.7904 0.7858 0.7930 0.8172
with S3 ranker only 0.2601 0.4901 0.6157 0.7880 0.8852 0.8037 0.7894 0.7849 0.7924 0.8168
OneRanker S2 0.2214 0.4257 0.5448 0.7291 0.8435 0.7566 0.7460 0.7440 0.7579 0.7876
w/o Target 0.2080 0.4030 0.5203 0.7066 0.8276 0.7375 0.7286 0.7285 0.7465 0.7786
w/o Target & MDA 0.2035 0.3922 0.5066 0.6949 0.8212 0.7360 0.7273 0.7275 0.7453 0.7776
Experimental results, as summarized in Table 2, demonstrate that each core module in our design contributes significantly to the overall performance. We primarily use HR@5 and NDCG@5 as the focal metrics for this ablation analysis to evaluate both hit accuracy and ranking quality.
In Step 2 , compared to the Base (OneRanker S2) performance (HR@5: 0.5448, NDCG@5: 0.7440), removing the Target information (w/o Target) causes a notable decline, with HR@5 dropping to 0.5203 (-4.5%) and NDCG@5 decreasing to 0.7285 (-2.1%), which strongly proves the necessity of the Fake-Item-Token mechanism, prompting the model to evolve from static user representations to dynamic-aware item semantic distribution representations, thereby capturing users’ real-time intentions more accurately. Furthermore, the simultaneous removal of Target information and the Multi-task Decoupled Architecture (w/o Target & MDA) leads to a further decline to 0.5066 in HR@5 and 0.7275 in NDCG@5, underscoring the critical role of the decoupled architecture in mitigating optimization tension within the shared representation space, effectively preventing interference between business value modeling and user interest modeling.
In Step 3, using the OneRanker with S3 ranker only (HR@5: 0.6157, NDCG@5: 0.7849) as the baseline, the model already demonstrates a leap-forward boost compared to S2-only models, proving that fine-grained target-aware mechanisms are fundamental to effective ranking. Building upon this, the introduction of S2 information injection (HR@5: 0.6161, NDCG@5: 0.7858) ensures that the ranker reuses the generation logic to better calibrate its predictions within the ”generation context”. Furthermore, incorporating the Distributional Consistency ${L}_{\text{DC}}$ (HR@5: 0.6173, NDCG@5: 0.7865) explicitly constrains the output distribution between stages, mitigating target shift through consistency regularization. Ultimately, the OneRanker (Full) model achieves the optimal HR@5 of 0.6213 and NDCG@5 of 0.7904. This improvement stems from a collaborative effect: while S2 provides enriched feature injection to S3, the Distributional Consistency loss facilitates essential value-driven feedback from S3 back to S2. This closed-loop interaction between input consistency and output consistency allows the model to reach a higher ceiling of performance.